### v2

@@ 3.2 Choice of hyperparameter values
Figure 1 illustrates how the algorithm proceeds if the user is interested in highly infrequent or highly frequent itemsets. A bottom-up search is being used in the first case; this bottom-up traversal of the tree allows for highly infrequent itemsets to be detected and any of their supersets to be removed from subsequent searching. For instance, in Figure 1(a) the itemset consisting of $\boldsymbol{X}_{1}=1$ and $\boldsymbol{X}_{2}=1$ is infrequent, thus the itemset that further includes the value of $\boldsymbol{X}_{3}$ (no matter what that may be) will not be considered in the search. If highly frequent itemsets are of interest instead, the method will proceed using a top-down approach instead, as shown in Figure 1(b). In this case, if the itemset consisting of $\boldsymbol{X}_{1}=1$ and $\boldsymbol{X}_{3}=1$ is flagged as frequent, all its subsets will also be marked as such and will therefore not be included in the rest of the search.
Our proposal for the score includes an additional hyperparameter, that is the exponent term $r>0$. This term controls the influence of itemsets of larger length in the computation of the score of nominal outlyingness. As [24] pointed out, an observation with frequent itemsets of unit length and $k\leq p$ infrequent itemsets of length two should not be seen as suspicious to be an outlier as an observation with $k$ infrequent itemsets of unit length. Our proposed framework extends this idea by letting the user decide the score contribution of itemsets of greater length. A larger $r$ value will yield a smaller augmentation of $s(\cdot)$ as $\lvert d\rvert$ increases, hence providing additional flexibility and control of how the score of outlyingness is being computed. We cannot recommend a specific value for $r$ as that is typically context-specific and depends on the weight the user wants to assign to highly infrequent (or frequent) itemsets of greater length. For instance, a reasonable value for $r$ would be $r\leq 3$; in case we set $r=3$, we would need $\sigma_{d}/\operatorname{supp}(d)>8$ for a two-dimensional itemset $d$ to contribute to $s(\cdot)$ as much as a specific outlying level on its own. While setting $r>3$ is certainly possible, we suggest that $r$ should not be much larger than the recommended value of $r=3$, so that longer sequences can also contribute to the score, at least to some extent.
@@ 3.3 Nominal outlyingness depth & Contribution matrix
We introduce the concept of nominal outlyingness depth, as an extension of our proposed method for quantifying nominal outlyingness. This is an additional metric that we introduce for meta-analysis purposes, so that the scores are easier for the user to understand and interpret. As we have illustrated in Figure 1, the way the score $s(\cdot)$ is being computed is a hierarchical process and can therefore be represented using a tree-based structure. This has motivated us to come up with a concept inspired from the continuous anomaly detection literature and more precisely from the Isolation Forest algorithm [33]. In the Isolation Forest algorithm, the outlyingness of each observation can be assessed by its average depth; that is the average number of random splits that it takes for the observation to be isolated and appear in a region that includes no other points. We proceed in a similar manner and introduce a Nominal Outlyingness Depth concept (NOD), which we denote by $\operatorname{NOD}(\cdot)$, given by:
$\operatorname{NOD}(\boldsymbol{x}_{i})=\left.\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{\operatorname{MAXLEN}}\end{subarray}}\lvert d\rvert\middle/\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{\operatorname{MAXLEN}}\end{subarray}}\right.1,$ (4)
@@
The rationale behind Expression (4) is that for each observation $\boldsymbol{x}_{i}$ we compute the average itemset length over all infrequent itemsets included in $\boldsymbol{x}_{i}$. A similar Expression to (4) can be computed if highly frequent itemsets are of interest; the only modification is changing the interval $(\sigma_{d},n]$ to $[0,\sigma_{d})$ and $\lvert d\rvert$ on the numerator becomes $\operatorname*{\operatorname{MAXLEN}}-\lvert d\rvert+1$, to account for the fact that we use a top-down search instead. One can also consider the definition of the nominal outlyingness depth using a tree representation; that would be the average traversal depth of highly infrequent or highly frequent itemsets, where the depth is relative to a base level that depends on whether a bottom-up or a top-down search is being used. Such an illustration is given in Figure 2 for the toy example of $\boldsymbol{x}=(1,1,1)^{\intercal}$. For instance, according to Figure 2(b), we are looking for highly infrequent itemsets and thus we use a bottom-up approach. The nominal outlyingness depth in this case is given by $(1+3)/2=2$. Similarly for the case of highly frequent itemsets, where a top-down search is being employed as in Figure 2(b), the nominal outlyingness depth would be $(2+3)/2=5/2$. The depth value ranges from a unit up to $\operatorname*{\operatorname{MAXLEN}}$, with a lower value corresponding to a larger degree of outlyingness, except for observations with no highly frequent or highly infrequent itemsets; these have a zero depth.
@@
The final concept we introduce is that of the contribution matrix, which we denote by $\boldsymbol{\mathrm{C}}$. This is a $(n\times p)$-dimensional matrix, with the $(i,j)$th entry being the contribution of variable $j$ to the nominal outlyingness score of $\boldsymbol{x}_{i}$. The contribution of the $j$th variable to the nominal score of outlyingness of the $i$th observation for highly infrequent itemsets is given by:
$c_{i,j}=\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ j\subseteq d\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{\operatorname{MAXLEN}}\end{subarray}}\frac{\sigma_{d}}{\operatorname{supp}(d)\times\lvert d\rvert^{r+1}},\quad r>0,$ (5)
@@
As a result, all the elements of the contribution matrix will be non-negative and the sum of each row is in fact the score of nominal outlyingness for the relevant observation, that is $\sum_{j=1}^{p}c_{i,j}=s(\boldsymbol{x}_{i})$. The main idea behind Expression (5) is that if an itemset is found to be highly infrequent, all variables involved will get an equal contribution to the augmentation of the score that is caused by this sequence of categorical levels. One can therefore use the contribution matrix to facilitate understanding of the source of nominal outlyingness of the data by seeing which variables are involved in highly infrequent (or frequent) itemsets, causing an increase of the score of nominal outlyingness.
@@ 4 Properties of the proposed score
We now explore some properties of our proposed score of nominal outlyingness. Notice that we focus exclusively on the case of an outlier being defined as consisting of highly infrequent itemsets; extending these properties to the frequent case is straightforward. We are mainly interested in the range of values the score can take, as well as on when it is maximised; essentially, we would like to know under what circumstances an observation is extremely likely to be outlying and how the associated hyperparameters influence its score.
@@
The maximum value of the nominal score of outlyingness $s(\boldsymbol{x}_{i})$ for infrequent itemsets in a data set with $p$ nominal variables is attained for all itemsets of length $1\leq k\leq p$ appearing just once.
@@
We assume that we have $p\geq 2$ nominal variables (for the case of $p=1$, the result follows immediately by the definition of the nominal score) and recall the formulation of the discrete score for an observation $1\leq i\leq n$ as given in Expression (1):
$s(\boldsymbol{x}_{i})=\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{\operatorname{MAXLEN}}\end{subarray}}\frac{\sigma_{d}}{\operatorname{supp}(d)\times\lvert d\rvert^{r}},\quad r>0.$
Assuming a large $\operatorname*{\operatorname{MAXLEN}}$ value (to avoid having many restrictions on the calculation of the score), it is easy to see that itemsets of unit support with a large minimum support threshold value $\sigma_{d}$ yield a higher increase in the score. This case corresponds to highly misspecified large probabilities for itemsets that only appear once in the data, for which $\sigma_{d}=n-\delta$, where $\delta$ is typically a small positive integer. Since $\sigma_{d}$ decreases as $\lvert d\rvert$ gets larger and as this is a parameter that is hard to determine in advance, we will stick with the maximisation of the $1/\lvert d\rvert^{r}$ term, assuming the maximum possible $\sigma_{d}$ value of $\sigma_{d}=n-1$. Therefore, the total score becomes equal to the following expression:
@@
Notice that the first two terms inside the curly brackets correspond to the contribution of the itemsets of unit length that appear once in the data set, while the third term corresponds to the contribution of all possible itemsets of greater length (up to length $p$) to the score. The coefficients $\alpha_{i}$ represent the number of nominal variables which are included in infrequent itemsets (of unit support) of length $i+1$. We further impose the restriction that the sum of the $\alpha_{i}$’s is at most equal to $p$, since any itemsets of length $i+1$ are defined based on $i+1$ variables, and we are restricted to $p$ nominal features. Moreover, we could potentially have no infrequent itemsets of length $i+1$ (in which case $\alpha_{i}=0$) but if we do have any, then these should be observed in at least $i+1$ nominal features. Despite the abuse of notation, we assume that $C^{0}_{i+1}=0$ for convenience.
Our goal is to find the set of values $\boldsymbol{\alpha}=\left(\alpha_{1},\dots,\alpha_{p-1}\right)$ that maximise $\mathcal{A}$ and more precisely, we aim to show that this is achieved either when all the $\alpha_{i}$’s are equal to zero or when we are on the boundary of the solution space, i.e. when all the $\alpha_{i}$’s sum to $p$ and all of them but one are exactly zero. The first case corresponds to maximising the nominal score for all $p$ nominal features of an observation being unique within the data set, while the second is interpreted as achieving the maximum nominal score possible when all itemsets of length $i+1$ that can be generated from $p$ nominal variables appear just once and there exist no itemsets of greater or lower length which are infrequent. In both these cases, we basically end up with the conclusion that the maximum score is attained for all itemsets of one specific length occurring once in the data set.
For $p<2^{r+1}+1$, we would rather have $\alpha_{i}=0\ \forall i$ if:
@@
The above condition can be equivalently written, by some straightforward algebraic manipulations, as:
@@
Since we assume that all the $\alpha_{i}$’s are non-negative, the expression above can be negative if and only if $\exists\alpha_{i}>0$ such that:
@@
where $\Gamma(\cdot)$ is the gamma function. If $F<0\ \forall\alpha_{i}>0$, we can ensure that $\mathcal{A}$ is maximised for all the $\alpha_{i}$’s being equal to zero. Since Expression (6) is maximised when $i=1$ and $\alpha_{1}=p$, we obtain:
@@
We now assume $p\geq 2^{r+1}+1$. We first show that getting to the solution boundary is always preferable when having one non-zero $\alpha_{i}$ and then, we show that setting $\alpha_{j}$ equal to a non-zero value, where $i\neq j$, decreases the value of $\mathcal{A}$. We begin by assuming that there exists an index $j$ such that $\alpha_{j}>0$ and $\alpha_{i}=0\ \forall i\neq j$ and we look at what happens to the value of $\mathcal{A}$ if we decrease $\alpha_{j}$ by a unit. Due to the restriction on the values that $\alpha_{j}$ can take, we need to assume that $j\neq p-1$, since $\alpha_{p-1}$ can only be equal to 0 or $p$. We denote the loss that results from decreasing $j$ by a unit by $G_{1}$ and that is equal to:
@@
When decreasing $\alpha_{j}$ from its minimum possible non-zero value of $j+1$ to zero, we define $G_{2}$ to be the loss corresponding to this decrease:
@@
Hence, for $G<0\ \forall j\in\mathbb{Z}^{+}$, which implies $\mathcal{A}$ increases when going from $\alpha_{j}=j+1$ to $\alpha_{j}=0$. Recalling that $\alpha_{j}>j$, we may write $\alpha_{j}=j+z$, where $z\in\{1,\ldots,p-j\}$. The expression for $G_{1}$ then becomes:
@@
The above expression is increasing in $z$ (this can be seen by considering the ratio $\Gamma(j+z)/\Gamma(z)$ and using Stirling’s approximation for the gamma function) and decreasing in $j$. Thus $G_{1}$ is minimised when $j=p-2$ and $z=1$, in which case:
@@
The final part of the proof consists of showing that if we are on the boundary of the solution space, with $\alpha_{j}=p$ and $\alpha_{i}=0\ \forall i\neq j$, then activating $\alpha_{k}$ so that $\alpha_{k}=k+1$ (where $k\neq j$) leads to a decrease in the value of $\mathcal{A}$. Now clearly, if $\alpha_{k}$ is set equal to $k+1$, the value of $\alpha_{j}$ has to drop to $p-(k+1)$, so that the boundary constraint is not violated. We define the loss incurred by activating $\alpha_{k}$ by $H$:
@@
×$\times\left\{\frac{\Gamma\left(p+1\right)}{\Gamma\left(p-j\right)}-\frac{\Gamma\left(p-k\right)}{\Gamma\left(p-k-j-1\right)}\right\}-\frac{1}{(k+1)^{r}}.$
We distinguish between two possible cases, namely $j<k$ and $j>k$. In the former case, we obtain:
@@
×$\times\left\{\frac{\Gamma\left(p+1\right)}{\Gamma\left(p-k+\beta\right)}-\frac{\Gamma\left(p-k\right)}{\Gamma\left(p-2k+\beta-1\right)}\right\}-$
@@
×$\times\left\{\frac{\Gamma\left(p+1\right)}{\Gamma\left(p-k-\beta\right)}-\frac{\Gamma\left(p-k\right)}{\Gamma\left(p-2k-\beta-1\right)}\right\}-$
@@
where $\beta\in\mathbb{Z}^{+}$. These expressions for $H$ give us additional restrictions, which are that $p-2k+\beta-1>0$ and $p-2k-\beta-1>0$, so that all terms are well-defined. We now show that $H>0\ \forall k\neq j$ just for the case of $j<k$; the second case follows analogously. Let the first term of $H$ be denoted by $H_{1}$ and denote the second one by $H_{2}$. It is obvious that $\beta=1$ minimises $H_{1}$. Thus, we have:
@@
where $Q(k)=\Gamma(p+1)/\Gamma(p-k)-\Gamma(p-k)/\Gamma(p-2k)$. It is also easy to see that $Q(k)$ is increasing for $k$, therefore its minimum is attained when $k=2$. In that case, we have:
@@
Thus, the minimum value of $H<1$ which shows $H>0\ \forall k\neq j$. As a result activating $\alpha_{k}$ is not preferred as it leads to a positive loss value.
@@
    If $p\geq 2^{r+1}+1$ and we are on the boundary of the solution space with just one parameter $\alpha_{j}$ being equal to $p$, then activating any other parameter $\alpha_{k}$ (where $j\neq k$) so that it is not longer equal to zero, leads to a lower value of $\mathcal{A}$.
@@
The maximum value of the score of nominal outlyingness $s(\boldsymbol{x}_{i})$ for the $i$th observation of a data set with $p$ nominal variables is given by $p(n-1)$ for $p<2^{r+1}+1$, and by the following expression for $p\geq 2^{r+1}+1$:
$s^{\mathrm{max}}(\boldsymbol{x}_{i})=(n-1)\frac{\dbinom{p}{\min\left\{\operatorname*{\operatorname{MAXLEN}},\left\lfloor\frac{p-r}{2}\right\rfloor\right\}}}{\left(\min\left\{\operatorname*{\operatorname{MAXLEN}},\left\lfloor\frac{p-r}{2}\right\rfloor\right\}\right)^{r}},$
[…]
@@
As proven in Proposition 1, if we assume that the largest possible value of $\sigma_{d}/\operatorname{supp}(d)$ is given by $n-1$, the score of nominal outlyingness $s$ is maximised when we only have highly infrequent (or frequent) itemsets of unit length, for $p<2^{r+1}+1$. In this case, it is obvious that the total number of such itemsets is equal to $p$, thus the score becomes equal to $(n-1)p$.
Now for the case of $p\geq 2^{r+1}+1$, as previously shown, the largest possible score can be achieved on the boundary of the solution space; that is for all itemsets of a specific length $k\leq p$ having unit support. The total score is then given by $(n-1)\times C^{p}_{k}/{k^{2}}$,where $C$ is the binomial coefficient. We therefore seek to find the itemset length $k^{*}\in\mathbb{Z}^{+}$ that maximises this expression of the score, for fixed $p$. This means that the score for $k=k^{*}-1$ and for $k^{*}+1$ (we assume $1<k^{*}<p$) will be less than that for $k^{*}$, or equivalently (omitting the common $(n-1)$ term):
@@
Starting with Expression (7), we have:

### v3

@@ 3.1 Score of nominal outlyingness
A bottom-up search is used when infrequent itemsets are of interest; this bottom-up traversal of the tree allows for highly infrequent itemsets to be detected and any of their supersets to be removed from subsequent searching. For instance, in Figure 1(a) the itemset consisting of $\boldsymbol{X}_{1}=1$ and $\boldsymbol{X}_{2}=1$ is infrequent, thus the itemset that further includes the value of $\boldsymbol{X}_{3}$ is not considered in the subsequent search. If highly frequent itemsets are of interest instead, the method proceeds analogously using a top-down approach instead, as shown in Figure 1(b). Therefore, assuming that $\operatorname{supp}((\boldsymbol{X}_{1},1),(\boldsymbol{X}_{2},1))=v$, $\operatorname*{L_{\text{max}}}>2$, and the corresponding minimum support threshold value is $u>v$, we get $s(\boldsymbol{x})=u/(2^{r}v)$; this is the same for any observation satisfying $\boldsymbol{X}_{1}=1$ and $\boldsymbol{X}_{2}=1$.
3.2 Choice of hyperparameter values
The hyperparameters involved in the calculation of the score $s(\cdot)$ are the minimum support threshold $\sigma_{d}$, the maximum length of sequences considered $\operatorname*{L_{\text{max}}}$, and the exponent term $r$. We start by looking at the minimum support threshold value $\sigma_{d}$. This is defined with respect to the assumed underlying Multinomial distribution for each itemset. Assume without loss of generality that we are working with itemsets $d$ consisting of the first $\lvert d\rvert$ nominal variables. For such itemsets, denoted as $\boldsymbol{X}_{[1:\lvert d\rvert]}$, the underlying Multinomial distribution that we consider is
$\boldsymbol{X}_{[1:\lvert d\rvert]}=\boldsymbol{x}\mid d\sim\text{Multinomial}\left(1,\boldsymbol{p}^{d}\vcentcolon=\bigotimes\limits_{i=1}^{\lvert d\rvert}\boldsymbol{\pi}_{i}\right),$
where $\otimes$ is the Kronecker product. Notice how this formulation implicitly assumes independence of the nominal variables; a dependence structure between the nominal variables does not change anything besides the definition of the probability vector, which has to be explicitly provided by the user instead.
In order to check if the support of an itemset is within a sensible range of values under the assumed model, we construct an upper one-sided $100(1-\alpha)\%$ confidence interval for the probability of observing itemset $d$. Since all itemsets of length $\lvert d\rvert$ form a Multinomial distribution with $\prod_{i=1}^{\lvert d\rvert}\ell_{i}$ possible outcomes, we use the method of [41] to construct $100(1-\alpha)\%$ simultaneous confidence intervals for the Multinomial proportions. The two-sided $100(1-\alpha)\%$ confidence intervals for the components of the Multinomial proportions vector $\boldsymbol{p}^{d}$ are given by
In a similar manner, the upper and lower one-sided $100(1-\alpha)\%$ confidence intervals are given by
Multiplying the lower bound of each upper one-sided interval by the sample size $n$ yields bounding values for the support of each itemset which are used as our values for $\sigma_{d}$; these are no longer universal but specific to each itemset and its assumed frequency.
As the itemset length grows, the associated contingency tables become more sparse. This leads to a serious issue if one assumes that $p_{i}^{d}\overset{\lvert d\rvert\rightarrow p}{\longrightarrow}0$, since almost any itemset is considered as highly infrequent and there is a substantial and potentially prohibitive increase in computational resources required. Determining a stopping criterion $\operatorname*{L_{\text{max}}}$ so that it reflects on our concerns regarding sparsity can be done by requiring all upper thresholds $\sigma_{d}$ to be at least equal to two for all itemsets $d$. The rationale behind this is that if $\sigma_{d}<2\ \forall d:\lvert d\rvert=M$, then only itemsets of length $M\leq p$ with unit support are not flagged as infrequent. This is a restrictive case leading to most (if not all) observations having their scores augmented due to being more frequent than expected. Thus, we define $\operatorname*{L_{\text{max}}}$ as follows
$\operatorname*{L_{\text{max}}}=\max\left\{\right$. $1\leq M\leq p:\sigma_{d}\geq 2$
The exponent term $r>0$ controls the influence of longer itemsets in the nominal outlyingness score. Based on the principle of [25] which states that an observation with frequent itemsets of unit length and $k\leq p$ infrequent itemsets of length two should not be seen as suspicious to be an outlier as an observation with $k$ infrequent itemsets of unit length, our framework allows the user to decide the score contribution of longer itemsets. The choice of $r$ is context-specific and may be selected based on the level of sparsity in the data set. In sparse datasets, where many longer itemsets occur rarely by construction, high values of $r$ may overly penalise these itemsets, while in denser data, moderate values allow longer patterns to contribute meaningfully. For practical purposes, we suggest restricting to $r\leq 3$; when $r=3$, a two-dimensional itemset requires $\sigma_{d}/\operatorname{supp}(d)>8$ to match the contribution of a single attribute. Larger values are possible, but we advise against setting $r$ too large, so that longer itemsets can still play a role without dominating the score.
@@ 3.3 Nominal outlyingness depth & Contribution matrix
We introduce the concept of nominal outlyingness depth, as an additional metric for meta-analysis purposes. As illustrated in Figure 1, the score $s(\cdot)$ is being computed in a hierarchical process and can therefore be represented using a tree-based structure. This has motivated us to come up with a concept inspired from the continuous anomaly detection literature and more precisely from the Isolation Forest algorithm [33], where outlyingness is assessed by the average depth; that is the average number of random splits that it takes for an observation to be isolated in the $p$-dimensional Euclidean space. We introduce a Nominal Outlyingness Depth concept, which we denote by $\operatorname{NOD}(\cdot)$, given by
$\operatorname{NOD}(\boldsymbol{x}_{i})=\left.\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{L_{\text{max}}}\end{subarray}}\lvert d\rvert\middle/\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{L_{\text{max}}}\end{subarray}}\right.1,$ (4)
@@
For each observation $\boldsymbol{x}_{i}$ we compute the average itemset length over all infrequent itemsets included in $\boldsymbol{x}_{i}$. A similar Expression to (4) can be formulated if highly frequent itemsets are of interest. Using a tree representation, NOD is defined as the average traversal depth of highly infrequent or highly frequent itemsets, where the depth is relative to a base level that depends on whether a bottom-up or a top-down search is being used. Such an illustration is given in Figure 2 for the toy example of $\boldsymbol{x}=(1,1,1)^{\intercal}$. For instance, according to Figure 2(a), we are looking for highly infrequent itemsets and thus we use a bottom-up approach. The nominal outlyingness depth in this case is given by $(1+2)/2=1.5$. Similarly for the case of highly frequent itemsets, $\text{NOD}(\boldsymbol{x})(2+3)/2=2.5$. The depth value ranges from a unit up to $\operatorname*{L_{\text{max}}}$, with a lower value corresponding to a larger degree of outlyingness, except for observations with no highly frequent or highly infrequent itemsets; these have a zero depth.
We remark that the proposed NOD parallels ideas of classical statistical depth functions [see 50, for a review] which measure centrality via distances or directional projections. Here, centrality is instead captured by the frequency structure of the underlying discrete distribution. Observations supported mainly by longer infrequent itemsets are considered more outlying, while those involving shorter, common itemsets are viewed as deeper.
@@
The final concept we introduce is that of the contribution matrix, which we denote by $\boldsymbol{\mathrm{C}}$. This is a $(n\times p)$-dimensional matrix, with the $(i,j)$th entry being the contribution of variable $j$ to the nominal outlyingness score of $\boldsymbol{x}_{i}$. For highly infrequent itemsets, this contribution is given by
$c_{i,j}=\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ j\subseteq d\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{L_{\text{max}}}\end{subarray}}\frac{\sigma_{d}}{\operatorname{supp}(d)\times\lvert d\rvert^{r+1}},\quad r>0,$ (5)
@@
All the elements of the contribution matrix are non-negative and the sum of each row is the score of nominal outlyingness for the relevant observation. The main idea behind Expression (5) is that if an itemset is found to be highly infrequent, all variables involved receive an equal contribution to the augmentation of the score that is caused by this sequence of categorical levels. The contribution matrix can facilitate understanding of the source of nominal outlyingness of the data by seeing which variables are involved in highly infrequent (or frequent) itemsets.
3.4 The SONO algorithm
Based on the previous subsections, we present the pseudo-code for our algorithm in Algorithm 1, which we call SONO (Scores Of Nominal Outlyingness). The code assumes highly infrequent itemsets are of interest.
Algorithm 1 The SONO Algorithm for highly infrequent itemsets
1:Input: Data set $\boldsymbol{X}$ with $n$ observations, $p$ nominal variables, exponent term $r>0$, significance level $\alpha>0$, expected proportions of variable levels $(\boldsymbol{\pi}^{\intercal}_{1},\ldots,\boldsymbol{\pi}^{\intercal}_{p})^{\intercal}$. (Optional: $L_{\max}$ value)
2:Output: Scores of nominal outlyingness $\{s(\boldsymbol{x}_{i})\}_{i=1}^{n}$
3:
4:Estimate $L_{\max}$ based on $(\boldsymbol{\pi}^{\intercal}_{1},\ldots,\boldsymbol{\pi}^{\intercal}_{p})^{\intercal}$ (unless user-provided)
5:Initialise list of infrequent itemsets $\mathcal{I}\leftarrow\varnothing$
6:Set $M\leftarrow 1$
7:while $M\leq L_{\max}$ do
8:  Record supports of all itemsets not in $\mathcal{I}$ of length $M$ in $\boldsymbol{X}$
9:  Construct simultaneous $100(1-\alpha)\%$ upper CIs for each itemset of length $|d|$ and determine $\sigma_{d}$ values
10:  if $\exists d\notin\mathcal{I}$ such that $\operatorname{supp}(d)<\sigma_{d}$ and $|d|=M$ then
11:   Augment $s(\boldsymbol{x}_{i})$ for all $\boldsymbol{x}_{i}\supseteq d$
12:   Flag $d$ and its supersets as infrequent; $\mathcal{I}\leftarrow\mathcal{I}\cup\{d^{\prime}:d^{\prime}\supseteq d\}$
13:  end if
14:end while
@@ 4 Properties of the proposed score
We now explore some properties of our proposed score of nominal outlyingness. Notice that we focus exclusively on the case of an outlier being defined as consisting of highly infrequent itemsets; extending these properties to the frequent case is straightforward. We are mainly interested in the range of values the score can take, as well as on when it is maximised; essentially, we would like to know under what circumstances an observation is extremely likely to be outlying and how the associated hyperparameters influence its score. Due to the complexity of the problem, the following statements rely on certain assumptions that represent rather extreme cases but they still provide useful intuition for interpreting the scores.
@@
Let $\operatorname*{L_{\text{max}}}=p$ and assume the minimum support threshold value is given by its maximum possible value so that $\sigma_{d}/\operatorname{supp}(d)=n-\delta$, where $\delta\rightarrow 1$ for all itemsets $d$. The maximum value of the nominal score of outlyingness $s(\boldsymbol{x}_{i})$ for infrequent itemsets in a data set with $p\geq 2$ nominal variables is attained for all itemsets of length $1\leq k\leq p$ appearing just once.
@@
Recalling the score in Expression (1) and given these assumptions, the total score is equivalently written as
@@
The coefficients $\alpha_{i}$ represent the number of nominal variables which are included in infrequent itemsets (of unit support) of length $i+1$. We further impose the restriction that the sum of the $\alpha_{i}$’s is at most equal to $p$, since any itemsets of length $i+1$ are defined based on $i+1$ variables, and we are restricted to $p$ nominal features. Moreover, we may have no infrequent itemsets of length $i+1$ (in which case $\alpha_{i}=0$) but if we do have any, then these should be observed in at least $i+1$ nominal features. Despite the abuse of notation, we assume that $C^{0}_{i+1}=0$ for convenience, with $C$ denoting the binomial coefficient.
For $p<2^{r+1}+1$, we would rather have $\alpha_{i}=0\ \forall i$ if
@@
Equivalently
@@
Since we assume that all the $\alpha_{i}$’s are non-negative, the expression above can be negative if and only if $\exists\alpha_{i}>0$ such that
@@
[…]
@@
We now assume $p\geq 2^{r+1}+1$. Assuming that there exists an index $j$ such that $\alpha_{j}>0$ and $\alpha_{i}=0\ \forall i\neq j$, we look at what happens to the value of $\mathcal{A}$ if we decrease $\alpha_{j}$ by a unit. Due to the restriction on the values that $\alpha_{j}$ can take, we need to assume that $j\neq p-1$, since $\alpha_{p-1}$ can only be equal to 0 or $p$. We denote the loss that results from decreasing $j$ by a unit by $G_{1}$ and that is equal to
@@
When decreasing $\alpha_{j}$ from its minimum possible non-zero value of $j+1$ to zero, we define $G_{2}$ to be the loss corresponding to this decrease
@@
Hence, $G_{2}<0\ \forall j\in\mathbb{Z}^{+}$Recalling that $\alpha_{j}>j$, we may write $\alpha_{j}=j+z$, where $z\in\{1,\ldots,p-j\}$. The expression for $G_{1}$ then becomes
@@
The above expression is increasing in $z$ and decreasing in $j$. Thus $G_{1}$ is minimised when j=p2j=p-2 and $z=1$, in which case
@@
The final part of the proof consists of showing that if we are on the boundary of the solution space, with $\alpha_{j}=p$ and $\alpha_{i}=0\ \forall i\neq j$, then activating $\alpha_{k}$ so that $\alpha_{k}=k+1$ (where $k\neq j$) leads to a decrease in the value of $\mathcal{A}$. Now clearly, if $\alpha_{k}$ is set equal to $k+1$, the value of $\alpha_{j}$ has to drop to p(k+1)p-(k+1), so that the boundary constraint is not violated. We define the loss incurred by activating $\alpha_{k}$ by $H$ as
@@
$\times\left\{\frac{\Gamma\left(p+1\right)}{\Gamma\left(p-j\right)}-\frac{\Gamma\left(p-k\right)}{\Gamma\left(p-k-j-1\right)}\right\}-\frac{1}{(k+1)^{r}}.$
In the case $j<k$, we obtain
@@
$\times\left\{\frac{\Gamma\left(p+1\right)}{\Gamma\left(p-k+\beta\right)}-\frac{\Gamma\left(p-k\right)}{\Gamma\left(p-2k+\beta-1\right)}\right\}$-
@@
$\times\left\{\frac{\Gamma\left(p+1\right)}{\Gamma\left(p-k-\beta\right)}-\frac{\Gamma\left(p-k\right)}{\Gamma\left(p-2k-\beta-1\right)}\right\}$-
@@
[…]
@@
where $Q(k)=\Gamma(p+1)/\Gamma(p-k)-\Gamma(p-k)/\Gamma(p-2k)$. It is also easy to see that $Q(k)$ is increasing for $k$, therefore its minimum is attained when $k=2$. In that case, we have
@@
[…]
@@
    If $p\geq 2^{r+1}+1$ and we are on the boundary of the solution space with just one $\alpha_{j}$ being equal to $p$, then activating any other parameter $\alpha_{k}$ (where $j\neq k$) so that it is not longer equal to zero, leads to a lower value of $\mathcal{A}$.
@@
Let $\operatorname*{L_{\text{max}}}=p$ and assume the minimum support threshold value is given by its maximum value so that $\sigma_{d}/\operatorname{supp}(d)=n-\delta$, where $\delta\rightarrow 1$ for all itemsets $d$. The maximum value of the score of nominal outlyingness $s(\boldsymbol{x}_{i})$ for the $i$th observation of a data set with $p$ nominal variables is given by p(n1)p(n-1) for $p<2^{r+1}+1$, and by the following expression for $p\geq 2^{r+1}+1$
$s^{\mathrm{max}}(\boldsymbol{x}_{i})=(n-1)\frac{\dbinom{p}{\min\left\{\operatorname*{L_{\text{max}}},\left\lfloor\frac{p-r}{2}\right\rfloor\right\}}}{\left(\min\left\{\operatorname*{L_{\text{max}}},\left\lfloor\frac{p-r}{2}\right\rfloor\right\}\right)^{r}},$
where $\lfloor\cdot\rfloor$ is the floor function (i.e. $\lfloor x\rfloor$ is the largest integer that is smaller than or equal to $x$).
@@
As proven in Proposition 1, under the given assumptions, the score of nominal outlyingness is maximised when we only have highly infrequent (or frequent) itemsets of unit length, for $p<2^{r+1}+1$. In this case, it is clear that the total number of such itemsets is equal to $p$, thus the score becomes equal to (n1)p(n-1)p.
For the case of $p\geq 2^{r+1}+1$, as previously shown, the largest possible score can be achieved on the boundary of the solution space; that is for all itemsets of a specific length $k\leq p$ having unit support. Therefore, we seek to find
$k^{*}=\operatorname*{arg\,max}\limits_{1\leq k\leq p}\frac{C^{p}_{k}}{k^{r}}.$
The following two conditions are thus obtained by considering $k=k^{*}-1$ and $k=k^{*}+1$
@@
Starting with Expression (7), we have